### v1

@@ Abstract
This paper studies the implementation of Bayes correlated equilibria in symmetric Bayesian nonatomic games, using direct information structures and obedient strategies. The main results demonstrate full implementation in a class of games with positive cost externalities. Specifically, if the game admits a strictly convex potential in every state, then for every Bayes correlated equilibrium outcome with finite support and rational action distributions, there exists a direct information structure that implements this outcome under all equilibria. When the potential is only weakly convex, we show that all equilibria implement the same expected social cost. Additionally, all Bayes correlated equilibria, including those with infinite support or irrational action distributions, are approximately implemented.
Keywords: Bayes correlated equilibrium, Bayesian Wardrop equilibrium, congestion games, information design, nonatomic games, potential games.
@@ 1 Introduction
In this paper, we study partial and full implementation through information design in symmetric cost games with incomplete information and a continuum of a priori uninformed players. A perfectly informed designer can choose how to correlate players’ information with the payoff-relevant state to influence their actions, for example, to minimize a cost function such as the social cost. In this context, the concept of Bayes correlated Wardrop equilibrium (BCWE) (the analog of a Bayes correlated equilibrium in games with a finite number of players) characterizes the set of state-dependent probability distributions over action distributions (hereafter called flows) that can be induced in an incentive-compatible way by some information structure. [^4] Koessler et al., (2024) show that the set of BCWE is the natural limit of the set of Bayes correlated equilibria (Bergemann and Morris,, 2016) in finite games with a large number of players. Furthermore, as shown in Koessler et al., (2024), regardless of the designer’s objective, a designer-optimal BCWE can be formulated as a solution to an optimization program with a bounded number of variables and constraints, determined by the number of actions and states.
@@
We show that approaching information design through the lens of nonatomic games provides important insights, particularly for achieving full implementation through direct recommendation systems. There is full implementation when all equilibria implement the same outcome for the chosen information structure. This property is crucial because it ensures the robustness of the information design problem to adversarial equilibrium selection. When only some equilibria induce the desired outcome, we talk about partial implementation.
@@
Our basic game is a nonatomic game with symmetric and incomplete information, thereby extending the class of population games studied by Sandholm, (2001, 2010) to the Bayesian setting. [^5] Sandholm, (2001, 2010) allow for multiple populations that differ in their preferences and sets of feasible actions. For notational simplicity, we assume in this paper that the basic game involves only one population, so players are ex-ante symmetric. We expect our results to hold more generally for multiple populations. We extend this basic game to asymmetric information by considering a family of simple information structures (as defined in Wu et al.,, 2021): The set of players is partitioned into a finite number of populations; within any given population, all players receive the same signal. Such an information structure is a direct information structure if the set of possible signals is identical to the set of possible actions. A Bayesian Wardrop equilibrium (BWE) (the analog of Bayes-Nash equilibrium for games with finite number of players) of the game extended with an information structure is said to be obedient if it is such that all players in all populations follow their recommended action.
@@
We first show that every BCWE with finite support and rational flows can be partially implemented as an obedient BWE with an appropriate direct information structure (Theorem 4.1). Hence, a version of the revelation principle (Myerson,, 1982) applies to symmetric nonatomic Bayesian games. Next, we turn to full implementation, which ensures that every BWE outcome of the extended game leads to the desired BCWE. We show that if the basic game admits a strictly convex potential in each state, then the BWE outcome is unique for every information structure (Theorem 5.3). Consequently, information design becomes extremely powerful in nonatomic games with a strictly convex potential because every BCWE with finite support and rational flows can be fully implemented—there exists a direct information structure such that the BCWE is implemented, regardless of the BWE selection.
@@
If in each state the basic game admits a potential that is only weakly convex, full implementation of a given outcome is usually not feasible via information design. Yet, we show that if the potential is weakly convex, then for every BCWE with finite support and rational flows, there exists a direct information structure such that the expected social cost is the same as the BCWE expected social cost regardless of the BWE (Theorem 5.7). In particular, the expected cost of the social cost-minimizing BCWE can be fully implemented.
@@
Our full implementation results apply, in particular, to congestion games with increasing cost functions. A congestion game has a strictly convex potential if the set of actions is equal to the set of resources and the cost of using a resource is strictly increasing in the number of users. When actions are general feasible subsets of resources and costs are added across resources, congestion games admit a weakly convex potential, provided that the cost of using each resource is weakly increasing in the number of users.
@@ 1.1 Related Literature
The notion of Wardrop equilibrium originated from the work of Wardrop, (1952) on traffic flow theory, where he formulated principles describing user-optimal and system-optimal routing in transportation networks. Beckmann et al., (1956) proved that a Wardrop equilibrium is the solution of a convex optimization problem. This idea was framed in the language of potential games (see, among others, Sandholm,, 2001, Roughgarden,, 2007, Roughgarden and Tardos,, 2007), using a generalization of the concept of potential game from the case of a finite number of players (Rosenthal,, 1973, Monderer and Shapley,, 1996) to the nonatomic case. A general framework for nonatomic potential games, extending beyond routing games, was studied by Sandholm, (2001), Hofbauer and Sandholm, (2009), Sandholm, (2010). Koessler et al., (2024) extend this framework to Bayesian games with nonatomic players by studying the notions of correlated and Bayes correlated Wardrop equilibrium. They showed that Bayes correlated Wardrop equilibria emerge as limits of action flows induced by Bayes correlated equilibria (Bergemann and Morris,, 2016) in finite-player games. Related nonatomic versions of correlated and Bayes correlated equilibria were studied by, e.g., Díaz et al., (2009), Tavafoghi and Teneketzis, (2018), and Zhu and Savla, (2022). Acemoglu et al., (2018) showed the relevance of information design in routing games, and proved that some classical paradoxes in routing games can be expressed in terms of negative value of information. Among the recent contributions on the topic, the reader is referred, for instance, to Massicot and Langbort, (2022), Cianfanelli et al., (2023), Ambrogio et al., (2024), Griesbach et al., (2024).
To the best of our knowledge, our full implementation results with direct information structures have no analog in the literature on information design, where even for binary action games, full implementation typically relies on complex indirect information structures, often with infinite type spaces. Existing tools and recent results in this area include, for example, Mathevet et al., (2020), Dworczak and Pavan, (2022), Hoshino, (2022), Inostroza and Pavan, (2023), Li et al., (2023), and Morris et al., (2024). The literature on information design is surveyed in Bergemann and Morris, (2019), Kamenica, (2019) and Forges, (2020). Dughmi, (2017) surveys the algorithmic literature on information design.
@@ 1.2 Organization of the Paper
The next section presents the model, including the basic game, its Bayes correlated Wardrop equilibria, information structures with finite populations and public signals within populations, and Bayesian Wardrop ($\varepsilon$-)equilibria of the basic game extended with an information structure. Section 3 provides two simple examples and illustrates how to implement some BCWE. Our partial and full implementation results are presented in Sections 4 and 5, respectively.
@@ 1.3 Notation
Given any finite set $A$, we define
@@
For the sake of simplicity, we use the shorthand $\Delta(A)$ to indicate $\Delta_{1}(A)$, which is just the simplex of probability measures on $A$. For a compact set $\mathcal{Y}$, $\Delta{(\mathcal{Y})}$ denotes the set of Borel probability measures on $\mathcal{Y}$.
@@ 2 The Model
We consider a basic game $\Gamma$, which is a symmetric Bayesian game with nonatomic players, a finite set of actions, and a finite set of states. Following Koessler et al., (2024) and most of the literature on nonatomic games, we adopt the convention that players minimize cost functions, which are simply the negatives of utility functions. Formally:
@@
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{c}\right)$ is given by a finite set of actions $\mathcal{A}$, a finite set of states $\Theta$, a full-support probability distribution over states $p\in\Delta(\Theta)$, and a profile of cost functions $\boldsymbol{c}=(c_{a})_{a\in\mathcal{A}}$, where for each $a\in\mathcal{A}$, $c_{a}\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ is a continuous cost function and $\mathcal{Y}=\Delta(\mathcal{A})$ is the set of flow profiles.
A flow profile (or action distribution) is denoted by $\boldsymbol{y}=(y_{a})_{a\in\mathcal{A}}\in\mathcal{Y}$, where $y_{a}$ is the flow on action $a$. For every $a\in\mathcal{A}$, $\theta\in\Theta$ and $\boldsymbol{y}\in\mathcal{Y}$, $c_{a}(\boldsymbol{y},\theta)$ is the cost of action $a$ when the state is $\theta$ and the flow profile in that state is $\boldsymbol{y}$. A transition probability $\mu\colon\Theta\to\Delta(\mathcal{Y})$ that associates a distribution over flows to each state is called an outcome of the game.
The notion of Bayes correlated Wardrop equilibrium in the next definition is the analog of the notion of Bayes correlated equilibrium (Bergemann and Morris,, 2016) in the context of nonatomic games (see Koessler et al.,, 2024).
@@
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}c_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\leq\sum_{\theta\in\Theta}p(\theta)\int y_{a}c_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta),\quad\forall\,a,b\in\mathcal{A}.$ (2.1)
In words, for each state $\theta\in\Theta$, a flow profile $\boldsymbol{y}$ is drawn at random according to the distribution $\mu(\,\cdot\mid\theta)$. The corresponding outcome is a BCWE if, for every action $a$ that has strictly positive flow in at least one state, conditional on being assigned $a$, the expected cost of playing $a$ is not strictly higher than the expected cost of playing any other action $b$. For example, suppose that for every $\theta$, $\boldsymbol{y}(\theta)$ is a Wardrop equilibrium of the complete information game $(\mathcal{A},(c_{a}(\,\cdot\,,\theta))_{a\in\mathcal{A}})$ i.e.,
$\forall a,b\in\mathcal{A},\ y_{a}(\theta)>0\implies c_{a}(\boldsymbol{y}(\theta),\theta)\leq c_{b}(\boldsymbol{y}(\theta),\theta).$ (2.2)
@@
As another example, if $\bar{\boldsymbol{y}}$ is a Wardrop equilibrium of the average game with complete information $(\mathcal{A},(\bar{c}_{a})_{a\in\mathcal{A}})$, with $\bar{c}_{a}(\,\cdot\,)=\sum_{\theta}p(\theta)c_{a}(\,\cdot\,,\theta)$, then $\mu(\theta)=\delta_{\bar{\boldsymbol{y}}}$ is a “non-revealing” BCWE. This corresponds to an equilibrium outcome of the basic game with no information.
We now introduce finite information structures and Bayesian Wardrop equilibria as defined in Wu et al., (2021).
@@
The game $\Gamma$ extended by information structure $\mathcal{I}$, denoted $(\Gamma,\mathcal{I})$, can be described as follows: for each state $\theta\in\Theta$, a type profile $\boldsymbol{\tau}\in\mathcal{T}$ is drawn with probability $\pi(\boldsymbol{\tau}\mid\theta)$ and, for every $k$, the entire population $k$ observes the type $\tau^{k}\in\mathcal{T}^{k}$. Then, actions are chosen.
@@
of interim flows profiles of all populations is called interim flow profile. For each profile of types $\boldsymbol{\tau}=(\tau^{k})_{k\in\mathcal{K}}$, we let $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ be the interim total flow on action $a$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}\in\Delta(\mathcal{A})$ be the interim total flow profile given $\boldsymbol{\tau}$. We let $\operatorname{\mathsf{P}}(\tau^{k})=\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)$ denote the total probability of type $\tau^{k}$ induced by $p$ and $\pi$, where $\mathcal{T}^{-k}\coloneqq\times_{j\neq k}\mathcal{T}^{j}$. Under the interim flow profile $\widehat{\boldsymbol{y}}$, the cost of population $k$, given type $\tau^{k}$ and action $a$ such that $y_{a}^{k}(\tau^{k})>0$, is
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\eqqcolon c_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (2.4)
The list $(c_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ is called an interim cost profile.
@@
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}c_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)+\varepsilon.$ (2.5)
That is, for each population $k$ that observes type $\tau^{k}$, a positive mass chooses action $a$ only if, up to $\varepsilon$, it has the smallest conditional expected cost. Notice that this is equivalent to the following inequality for all $k\in\mathcal{K}$, all $\tau^{k}\in\mathcal{T}^{k}$ and all $a,b\in\mathcal{A}$ such that $y_{a}^{k}(\tau^{k})>0$:
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)c_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)+\varepsilon\operatorname{\mathsf{P}}(\tau^{k}).$ (2.6)

### v2

@@ Abstract
This paper studies the implementation of Bayes correlated equilibria in symmetric Bayesian games with nonatomic players, using direct information structures and obedient strategies. The main results demonstrate full implementation in a class of games with negative payoff externalities, such as congestion and Cournot games. Specifically, if the game admits a strictly concave potential in every state, then for every Bayes correlated equilibrium outcome with finite support and rational action distributions, there exists a direct information structure that implements this outcome under all equilibria. When the potential is weakly concave, we show that all equilibria implement the same expected total payoff. Additionally, all Bayes correlated equilibria, including those with infinite support or irrational action distributions, are approximately implemented.
Keywords: Bayes correlated equilibrium, congestion games, aggregative games, information design, nonatomic games, population games, potential games.
@@ 1 Introduction
The fundamental equilibrium concepts used in information design for incomplete-information games with finitely many players, are Bayes Nash equilibrium and Bayes correlated equilibrium. The use of these concepts in games with a large number of players is often cumbersome and computationally intractable. Yet, large games with symmetries can be approximated by nonatomic games, whose natural equilibrium concept in the case of complete information—Wardrop equilibrium—is based on distributions over actions (called flows), rather than on the behavior of single players. This technique was recently extended to incomplete-information nonatomic games, and the concepts of Bayesian Wardrop equilibrium (BWE) and Bayes correlated Wardrop equilibrium (BCWE) were introduced and analyzed (see Koessler et al.,, 2025).
@@
In this paper, we study partial and full implementation through information design in symmetric nonatomic games with incomplete information. In these, there is a payoff-relevant state, to which information will be correlated by a perfectly informed designer whose goal is—for instance—to maximize total welfare. In this context, the concept of BCWE characterizes the set of state-dependent probability distributions over flows that can be induced in an incentive-compatible way by some information structure. Furthermore, regardless of the designer’s objective, a designer-optimal BCWE can be formulated as a solution to an optimization program with a bounded number of variables and constraints, determined by the number of actions and states.
@@
We show that approaching information design through the lens of nonatomic games provides important insights, particularly for achieving full implementation through direct recommendation systems. There is full implementation when all equilibria implement the same outcome for the chosen information structure. This property is crucial because it ensures robustness of the information design problem to adversarial equilibrium selection. When only some equilibria induce the desired outcome, we talk about partial implementation.
@@
Our basic game is a nonatomic game with symmetric and incomplete information. Our analysis extends the class of population games studied by Sandholm, (2001, 2010) to the Bayesian setting. [^4] Sandholm, (2001, 2010) allowed for multiple populations with different preferences and sets of feasible actions. For notational simplicity, in this paper we assume that the basic game involves only one population, so players are ex-ante symmetric. At the cost of cumbersome notation, our results can be generalized to multiple populations. We extend this basic game to asymmetric information by considering a family of simple information structures (as defined in Wu et al.,, 2021): The set of players is partitioned into a finite number of populations; within any given population, all players receive the same signal. Such an information structure is a direct information structure if the set of possible signals is identical to the set of possible actions. A BWE of the game extended with an information structure is said to be obedient if all players in all populations follow their recommended action.
@@
We first show that every BCWE with finite support and rational flows can be partially implemented as an obedient BWE with an appropriate direct information structure (Theorem˜4.1). Hence, a version of the revelation principle (Myerson,, 1982) applies to symmetric nonatomic Bayesian games. Next, we turn to full implementation, which ensures that every BWE outcome of the extended game leads to the desired BCWE. We show that if the basic game admits a strictly concave potential in each state, then the BWE outcome is unique for every information structure (Theorem˜5.5). Consequently, information design becomes very effective in nonatomic games with a strictly concave potential because every BCWE with finite support and rational flows can be fully implemented, i.e., there exists a direct information structure such that the BCWE is implemented, regardless of the BWE selection.
@@
If in each state the basic game admits a potential that is just weakly concave, full implementation of a given outcome may not be achieved, yet we get full payoff implementation. We show that for every BCWE with finite support and rational flows, there exists a direct information structure such that expected total payoff is the same as the BCWE expected total payoff regardless of the BWE (Theorem˜5.9). In particular, the expected payoff of the total payoff-maximizing BCWE can be fully implemented.
@@
Our full (payoff) implementation results apply, in particular, to congestion games with decreasing payoff functions and to Cournot games. A congestion game has a strictly concave potential if the set of actions is equal to the set of resources and the payoff of using a resource is strictly decreasing in the number of users. When actions are general feasible subsets of resources and payoffs are added across resources, congestion games admit a concave potential, provided that the payoff of using each resource is weakly decreasing in the number of users. A Cournot game admits a concave potential if inverse demand is strictly decreasing, and the potential is strictly concave when possible quantities for each firm are binary (e.g., entry and no entry).
@@ 1.1 Related Literature
The notion of Wardrop equilibrium originated from the work of Wardrop, (1952) on traffic flow theory, where he formulated principles describing user-optimal and system-optimal routing in transportation networks. Beckmann et al., (1956) proved that a Wardrop equilibrium is the solution of a convex optimization problem. This idea was framed in the language of potential games (see, among others, Sandholm,, 2001, Roughgarden,, 2007, Roughgarden and Tardos,, 2007), generalizing the case of a finite number of players (Rosenthal,, 1973, Monderer and Shapley,, 1996) to the nonatomic case. A general framework for nonatomic potential games, extending beyond routing games, was studied by Sandholm, (2001), Hofbauer and Sandholm, (2009), Sandholm, (2010). The class of aggregative games—subclass of potential games—was implicitly introduced by Dubey et al., (1980) and formalized by Corchón, (1994). Lahkar, (2017) studied this class in games with nonatomic players and a finite set of actions.
Koessler et al., (2025) extended this framework to nonatomic Bayesian games by studying the notions of correlated and Bayes correlated Wardrop equilibria. They showed that Bayes correlated Wardrop equilibria emerge as limits of action flows induced by Bayes correlated equilibria (Bergemann and Morris,, 2016) in finite-player games. [^5] In Koessler et al., (2025) players minimize cost functions, whereas here players maximize payoff functions, which are simply the negatives of cost functions. Related nonatomic versions of correlated and Bayes correlated equilibria were studied, e.g., by Díaz et al., (2009), Tavafoghi and Teneketzis, (2018), and Zhu and Savla, (2022). Acemoglu et al., (2018) showed the relevance of information design in routing games, and proved that some classical paradoxes in routing games can be expressed in terms of negative value of information. Among the recent contributions on the topic, the reader is referred, for instance, to Massicot and Langbort, (2022), Cianfanelli et al., (2023), Ambrogio et al., (2024), and Griesbach et al., (2024). In a concurrent work focusing on congestion games and using distributional approach, Massicot, (2025, chapter 5) considers infinite information structures and obtains implementation results related to ours.
To the best of our knowledge, our full implementation results with direct information structures have no analog in the literature on information design, where—even for binary action games—full implementation typically relies on complex indirect information structures, often with infinite type spaces. Existing tools and recent results in this area include, for example, Mathevet et al., (2020), Dworczak and Pavan, (2022), Hoshino, (2022), Inostroza and Pavan, (2025), Li et al., (2023), and Morris et al., (2024). The literature on information design is surveyed in Bergemann and Morris, (2019), Kamenica, (2019) and Forges, (2020). Dughmi, (2017) surveys the algorithmic literature on information design.
@@ 1.2 Organization of the Paper
Section˜2 presents the model, including the basic game, its Bayes correlated Wardrop equilibria, information structures with finite populations and public signals within populations, and Bayesian Wardrop ($\varepsilon$-)equilibria of the basic game extended with an information structure. Section˜3 provides two simple examples and illustrates how to implement welfare-maximizing BCWE. Our partial and full implementation results are presented in Sections˜4 and 5, respectively. In the Appendix, Section˜6 identifies a class of games for which deterministic flows maximize the designer’s payoff function. Appendix˜A contains a list of symbols and acronyms used throughout the paper.
@@ 1.3 Notation
Given any finite set $A$ and $\gamma>0$, we define
@@
For the sake of simplicity, we use the shorthand $\Delta(A)$ to indicate $\Delta_{1}(A)$, which is just the simplex of probability measures on $A$. For a compact set $\mathcal{Y}$, $\Delta{(\mathcal{Y})}$ denotes the set of Borel probability measures on $\mathcal{Y}$. For $h\in\mathbb{N}$, the symbol $C^{h}$ denotes the class of functions having an $h$-th derivative that is continuous in its domain.
@@ 2 The Model
We consider a basic game $\Gamma$, which is a symmetric nonatomic Bayesian game with a finite set of actions, and a finite set of states. Without any loss of generality, throughout the paper, the total mass of players is assumed to be $1$. To define the basic game, we need the concept of flow profile, which is the way the mass of players is distributed over the possible actions. Given a finite set of actions $\mathcal{A}$, a flow profile $\boldsymbol{y}\coloneqq(y_{a})_{a\in\mathcal{A}}\in\mathcal{Y}$, is a nonnegative vector in $\mathcal{Y}\coloneqq\Delta(\mathcal{A})$, and $y_{a}$ is the flow on action $a$.
@@
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{u}\right)$ is given by a finite set of actions $\mathcal{A}$, a finite set of states $\Theta$, a full-support probability distribution $p\in\Delta(\Theta)$ over states, and a profile of payoff functions $\boldsymbol{u}=(u_{a})_{a\in\mathcal{A}}$, where, for each $a\in\mathcal{A}$, $u_{a}\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ is continuous.
For every $a\in\mathcal{A}$, $\theta\in\Theta$ and $\boldsymbol{y}\in\mathcal{Y}$, $u_{a}(\boldsymbol{y},\theta)$ is the payoff of action $a$ when the state is $\theta$ and the flow profile in that state is $\boldsymbol{y}$. A transition probability $\mu\colon\Theta\to\Delta(\mathcal{Y})$ that associates a distribution over flows to each state is called an outcome of the game.
The notion of Bayes correlated Wardrop equilibrium in the next definition is the analog of the notion of Bayes correlated equilibrium (Bergemann and Morris,, 2016) in the context of nonatomic games (see Koessler et al.,, 2025).
@@
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\geq\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta),\quad\forall\,a,b\in\mathcal{A}.$ (2.1)
For each state $\theta\in\Theta$, a flow profile $\boldsymbol{y}$ is drawn at random according to the distribution $\mu(\,\cdot\mid\theta)$. If we divide the two sides of (2.1) by the ex-ante expected mass of players choosing $a$, then the left-hand-side is proportional to the expected payoff of playing $a$, conditionally of being recommended $a$, and the right-hand side is proportional to the expected payoff of playing $b$, conditionally of being recommended $a$. The corresponding outcome is a BCWE if, for every action $a$ that has strictly positive flow in at least one state, conditionally on being assigned $a$, the expected payoff of playing $a$ is greater or equal to the expected payoff of playing any other action $b$.
For example, suppose that for every $\theta$, $\boldsymbol{y}(\theta)$ is a Wardrop equilibrium of the complete information game $(\mathcal{A},(u_{a}(\,\cdot\,,\theta))_{a\in\mathcal{A}})$, i.e.,
$\forall a,b\in\mathcal{A},\ y_{a}(\theta)>0\implies u_{a}(\boldsymbol{y}(\theta),\theta)\geq u_{b}(\boldsymbol{y}(\theta),\theta).$ (2.2)
@@
As another example, if $\bar{\boldsymbol{y}}$ is a Wardrop equilibrium of the average game with complete information $(\mathcal{A},(\bar{u}_{a})_{a\in\mathcal{A}})$, with $\bar{u}_{a}(\,\cdot\,)=\sum_{\theta}p(\theta)u_{a}(\,\cdot\,,\theta)$, then $\mu(\theta)=\delta_{\bar{\boldsymbol{y}}}$ is a “non-revealing” BCWE. This corresponds to an equilibrium outcome of the basic game with no information.
We now consider finite information structures and Bayesian Wardrop equilibria as defined in Wu et al., (2021). Using finite information structure has the following merits: On one hand this avoids measurability issues related to private signals for uncountable sets of players. On the other hand, the approach is flexible enough to guarantee the possibility of arbitrarily fine information diversification.
@@
The game $\Gamma$ extended by information structure $\mathcal{I}$, denoted by $(\Gamma,\mathcal{I})$, can be described as follows: for each state $\theta\in\Theta$, a type profile $\boldsymbol{\tau}\in\mathcal{T}$ is drawn with probability $\pi(\boldsymbol{\tau}\mid\theta)$ and, for every $k$, the entire population $k$ observes the type $\tau^{k}\in\mathcal{T}^{k}$. Then, actions are chosen.
@@
of interim flows profiles of all populations is called interim flow profile. For each type profile $\boldsymbol{\tau}=(\tau^{k})_{k\in\mathcal{K}}$, we let $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ be the interim total flow on action $a$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}\in\Delta(\mathcal{A})$ be the interim total flow profile, given $\boldsymbol{\tau}$. We let $\operatorname{\mathsf{P}}(\tau^{k})=\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)$ denote the total probability of type $\tau^{k}$ induced by $p$ and $\pi$, where $\mathcal{T}^{-k}\coloneqq\times_{j\neq k}\mathcal{T}^{j}$. Under the interim flow profile $\widehat{\boldsymbol{y}}$, the payoff of population $k$, given type $\tau^{k}$ and action $a$, such that $y_{a}^{k}(\tau^{k})>0$, is
[…]
The list $(u_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ is called an interim payoff profile.
@@
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}u_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)-\varepsilon.$ (2.5)
That is, for each population $k$ that observes type $\tau^{k}$, a positive mass chooses action $a$ only if, up to $\varepsilon$, it yields the highest conditional expected payoff. Notice that this is equivalent to the following inequality for all $k\in\mathcal{K}$, all $\tau^{k}\in\mathcal{T}^{k}$ and all $a,b\in\mathcal{A}$ such that $y_{a}^{k}(\tau^{k})>0$:
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)u_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)-\varepsilon\operatorname{\mathsf{P}}(\tau^{k}).$ (2.6)